# Evaluating Language Models on Entity Disambiguation in Tables

### v1

@@ 5.2. Ablation studies
Figure 4. Accuracies achieved by TableLlama and TURL w.r.t. the number of candidates.
@@
To explain the drop in performances observed for TURL and TableLlama on the TURL-2K-red-LamAPI dataset, we have measured the accuracy w.r.t. the number of candidates per mention, with the intuition that the higher the number of candidates the lower the performances. Figure 4 reports the accuracy of TURL and TableLlama given a different number of candidates per mention, considering also the correct one. Our intuitions are empirically confirmed by observing a drop in performance for both approaches, with a more severe one for TURL. This could happen because TURL aggregates the candidates of every mention in a table and passes them through a Transformer, increasing the context length to at most $O(NK)$, where $N$ is the number of mentions in a table and $K$ is the maximum number of candidates retrieved per mention.
We ran an additional ablation study to measure the impact of the table’s metadata (e.g., the title of the Wikipedia page the table is found in, the section title or the table caption to name a few) on the final accuracy achieved by both TURL and TableLlama, testing both on the TURL-2K-red-LamAPI dataset, the only dataset with table metadata, with i) No-Meta, i.e., a setting where the page title, section title and the table caption are removed; ii) No-Header, i.e., the table header is changed to $[\text{col}0,\text{col}1,...,\text{col}N]$ and iii) No-Meta-No-Header, i.e., the combination of both i) and ii).
Figure 5. Accuracies achieved by TableLlama and TURL w.r.t. the presence of table’s metadata.
@@
From Figure 5, we observe that TURL is heavily dependent on the table metadata, with a severe drop in performance when both metadata and header are removed. On the other hand, TableLlama is almost unaffected by removing metadata from the prompt, indicating a higher generalisation capability and a greater focus on the context provided by the table itself rather than by the metadata).
5.3. Discussion
@@
Generative GTUM approaches with a high number of parameters seem to have interesting properties in terms of accuracy, generalisation, and robustness to the number of candidates that are processed and the table metadata, especially on MOOD and OOD data; however, specific STI approaches trained on in-domain data, and optionally fine-tuned on MOOD data, using a tiny fraction of these parameters can still outperform generalistic approaches with speed higher by an order of magnitude; these results may suggest that generative GTUM approaches are at the moment the best choice for processing small tables, while more specific entity disambiguation approaches may still be a better fit for applications on large business data. Generative approaches like TableLlama seem more promising than encoder-based ones in terms of performance, with a negligible risk of hallucinations; however, the comparison considered models of uncomparable size (7B vs 300M for TableLlama and TURL resp.), with TURL being the fastest approach among those that were tested. Regarding scalability, long context allows the encapsulation of large tables with a high number of candidates, but some tables cannot fit in the context, and when they do, the computation time and occupied memory increase proportionally to the table size: considering a small context table (few rows above and below the row containing the mention to be linked) to be fed to the model could be effective. On the budget side, training of models of the size of TableLlama has still enormous costs (48 A100-80GB for 9 training days), so devising generative methods based on smaller LLMs, e.g., Phi (abdin2024phi3, ), could be an interesting research direction, although more sophisticated approaches may be needed to achieve the same level of generalisation and reliability. Furthermore, TableLlama and TURL are not NIL aware and TableLlama, in particular, cannot return confidence scores associated with cell annotations, which may be useful when using these approaches to support revision: devising methods to estimate the uncertainty of labels computed by generative models may be an interesting research direction.
@@ 6. Conclusions and Future Works
STI, the process of annotating tabular data with information from background KGs, is proposed to support the understanding, interpretation and labeling of tables. Among the STI’s tasks, CEA, i.e.,  matching cell values to entities in the KG, is particularly relevant to support additional downstream transformation, integration, and enrichment processes, and is particularly subject to scalability constraints, e.g., when applied to tables with a large number of rows. LLMs pretrained with a vast amount of data have been applied to STI and CEA, complementing previous approaches based on heuristic and featured-based ML approaches. However, these different families of approaches have not been exhaustively examined on a common ground. In this work, we tackled this gap by selecting four representative approaches and comparatively evaluating them in terms of accuracy, generalisability, time, and memory requirements to better study their strengths and limitations, as well as their potential applications to different scenarios. We defined different evaluation settings, i.e., “in-domain”, “out-of-domain” and “moderately out-of-domain” for a better analysis of generalisability.
Our experiments suggest that an approach like TableLlama, based on a large generative LLM excels in accuracy and generalisation, as demonstrated by the results on MOOD and OOD data (see Table 4), at the price of an excessive execution and training time. TURL, an encoder-only model based on TinyBERT, is the most efficient, but lacks on generalisation capabilities: however, fine-tuning using data from similar distributions can lead to improvements with the new data at the price of a drop with pre-train data. Evaluating the impact of a larger encoder-based LLM on an approach like TURL could be an interesting research direction. However, our experiments suggest that specific STI models like Alligator, despite their limited number of parameters can still outperform generalistic models in IN-domain settings with a huge gain in efficiency.
Future works include the possibility to train a smaller and cheaper LLM-based model, e.g., Phi (abdin2024phi3, ), while enabling also the handling of NIL entities and a score associated with cell annotations. Another possible direction is to let both TURL and TableLlama be cell-based instead of table-based to reduce the occupied memory by both approaches and to enable standard augmentation techniques for the former.
The datasets, procedure source code for their creation, and evaluation code will be publicly accessible upon acceptance.

### v2

@@ 5.2. Ablation studies
Figure 4. Accuracies achieved by Alligator, GPT4o, TableLlama and TURL w.r.t. the number of candidates on the TURL-2K-red-LamAPI dataset.
To explain the drop in performances observed for TURL and TableLlama on the TURL-2K-red-LamAPI dataset, we have measured the accuracy w.r.t. the number of candidates per mention, with the intuition that the higher the number of candidates the lower the performances. Figure 4 reports the accuracy of TURL, TableLlama, GPT-4o and GPT-4o-mini given a different number of candidates per mention, also considering the correct one. Our intuitions are empirically confirmed by observing a drop in performance for both approaches, with a more severe one for TURL. This could happen because TURL aggregates the candidates of every mention in a table and passes them through a Transformer, increasing the context length to at most $O(NK)$, where $N$ is the number of mentions in a table and $K$ is the maximum number of candidates retrieved per mention. We ran an additional ablation study to measure the impact of the table’s metadata (e.g., the title of the Wikipedia page the table is found in, the section title, or the table caption to name a few) on the final accuracy achieved by TURL, TableLlama, GPT-4o, and GPT-4o-mini, testing on the TURL-2K-red-LamAPI dataset. We ran experiments with i) No-Meta, i.e., a setting where the page title, section title, and the table caption are removed; ii) No-Header, i.e., the table header is changed to $[\text{col}0,\text{col}1,...,\text{col}N]$ and iii) No-Meta-No-Header, i.e., the combination of both i) and ii). From Figure 5, we observe that TURL is heavily dependent on the table metadata, with a severe drop in performance when both metadata and header are removed. On the other hand, TableLlama and GPT-4o are almost unaffected by removing metadata from the prompt, indicating a higher generalisation capability and a greater focus on the context provided by the table itself rather than by the metadata.
5.3. Discussion
@@
Figure 5. Accuracies achieved by GPT4o, TableLlama and TURL w.r.t. the presence of table’s metadata on the TURL-2K-red-LamAPI dataset.
@@
Generative GTUM approaches with a high number of parameters have interesting properties in terms of accuracy, generalisation, and robustness to the number of candidates that are processed and the table metadata, especially on MOOD and OOD data, as found in Sections 5.1 and 5.2; however, specific STI approaches trained on in-domain data, and optionally fine-tuned on moderately-out-of-domain data, using a tiny fraction of these parameters can still be on par with generalistic approaches while being faster and occupying less memory. Generative approaches like GPT-4o and TableLlama are more promising than encoder-based ones in terms of performance, with a negligible risk of hallucinations; however, the comparison considered models of uncomparable size (7B vs 300M for TableLlama and TURL resp.), with TURL being the fastest approach among those that were tested. Regarding scalability, long context allows the encapsulation of large tables with a high number of candidates, but some tables cannot fit in the context, and when they do, the computation time and occupied memory increase proportionally to the table size: considering a small context table (few rows above and below the row containing the mention to be linked) to be fed to the model could be effective. On the budget side, training models of the size of TableLlama has still enormous costs (48 A100-80GB for 9 training days), so devising generative methods based on smaller LLMs, e.g., Phi (abdin2024phi3, ), could be an interesting research direction, although more sophisticated approaches may be needed to achieve the same level of generalisation and reliability. Moreover, Figure 6 shows the accuracy achieved by the tested approaches, both before and after MOOD fine-tuning, with respect to the overall price [^12] For TableLlama, TURL and Alligator, prices are estimated using the Azure Pricing Calculator, selecting the NC24ads_a100_v4 for running GPU inference with TableLlama and TURL, and D48ds_v5 for running CPU inference and hosting Alligator. GPT-4o and GPT-4o-mini prices are estimated from the OpenAI prices page, considering a token as being equal to 4 characters. paid to host machines to run inference on all the test data. The figure shows how GPT-4o, while being the best-performing model on average, is also the most expensive one; GPT-4o-mini, TableLlama, TableLlama-FT, and Alligator-FT are comparable in terms of accuracy and price paid. These results suggest that generative GTUM approaches are best suited when the tables to be annotated contain rich, highly semantic textual data. GPT-4o is the most accurate among these methods but also the most expensive. GPT-4o-mini and TableLlama offer a fair compromise between accuracy, cost, and computational demand. On the other hand, Alligator is more appropriate when signals from the knowledge graph (KG) have a significant impact on the final accuracy, especially when time and budget are limited. When fine-tuned, Alligator performs comparably to GPT-4o-mini and TableLlama in terms of accuracy and cost.
[…]
Figure 6. Per-approach accuracy over price. The price is computed given the overall time spent (in minutes) to run the evaluation over all the test data (15205 mentions).
Finally, TableLlama and TURL are not NIL aware and TableLlama and GPT-4o, in particular, cannot return confidence scores associated with cell annotations, which may be useful when using these approaches to support revision: devising methods to estimate the uncertainty of labels computed by generative models may be an interesting research direction.
@@ 6. Conclusions and Future Works
In this study, we addressed the lack of comprehensive comparative analyses among different approaches to Semantic Table Interpretation (STI), focusing specifically on the Cell Entity Annotation (CEA) task. We selected four representative methods—including large generative Large Language Models (LLMs) and specialized STI models—and evaluated them under common conditions. Our evaluation spanned multiple dimensions: accuracy, generalisation, execution time, memory usage, and cost. We tested these methods across various scenarios, i.e., “in-domain”, “out-of-domain” and “moderately out-of-domain”. Our experiments suggest that approaches based on a large generative LLM excels in accuracy and generalisation, as demonstrated by the results on MOOD and OOD data (cfr. Table 5 and Figure 3), at the price of an increased execution and training time. TURL, an encoder-only model based on TinyBERT, is the most efficient, but lacks on generalisation capabilities. Moreover, our experiments suggest that specific STI models like Alligator, despite their limited number of parameters, can still be a viable option when resources and budget are limited, with performance improvements obtained after fine-tuning. Finally, GPT-4o, while being the most accurate, is also the most expensive, with GPT-4o-mini that represents a better trade-off with respect to accuracy and cost. Future works include the possibility to train a smaller and cheaper LLM-based table model, e.g., Phi (abdin2024phi3, ), while enabling also the handling of NIL entities and a score associated with cell annotations. Additionally, we aim to augment LLM-based approaches with advanced prompting techniques like In-Context Learning or Chain-of-Thought reasoning. This enhancement could leverage the inherent reasoning capabilities of LLMs and facilitate human-in-the-loop interactions to guide the model’s behavior. Another potential direction is to modify models like TURL and TableLlama to operate at the cell level rather than the table level. This change could reduce memory consumption and allow for the application of standard data augmentation techniques, potentially improving both performance and efficiency.